PartnerObserving Log for 2007 Oct 07-08

Observers: J. Hill, V. Testa
Instrument Support: R. Bertram
Telescope Operator: A. Ceranski

00:49 - 25 bias frames

01:05 - 5 bias

01:10 - Open Dome. T=5degC D=-29degC (Hammer used by Vincenzo on the hygrometer was very effective!) Wind 4 m/sec from NE. Sky is beautifully clear except for a few cirrus moving away in the SE. Vincenzo says not quite photometric.

Skyflats

01:14 - SkyFlatTestU: 011823 counts/sec = 19000

01:18 - SkyFlatU_5 pos.angle=0, exp.time=3.0s
011943: counts = 45000
012025: " = 38000
012113: " = 31000

01:24 - Images are not arriving -- Telescopes stops tracking during skyflats. Breaker for HBS pump #1 has tripped off again. Reset breaker, but it trips again in 15 minutes. We used the local switch to disable pump #1, and now oil is happy. Lost 30 minutes to HBS.

01:26 - after 5 minutes timeout got another exposure
012640: counts = 7000
Need other 5 min for the next timeout. Too dark for U. Switch to g: SkyFlatTestg, although it might too dark for g-SLOAN too....
013030: counts = 3500

01:31 - SkyFlatTestG: 013143 counts/sec = 6500
01:32 - SkyFlatG_5; pos. angle=0; exp.time = 8.0s
013252: counts = 46000
013339: " = 37000
013426: " = 34000
013515: " = 29000
013603: " = 24000

We tried to reenable Actuator 140 during twilight, but it is solidly inoperable. This panic'd the mirror, so we lost 10 minutes of twilight to PMC-SX.

Had lots of problems with LBCFPIA not finding pupils during the initial active optics. It seems that the astigmatism created by Actuator 140 being off is making the pupil finding much more difficult. One edge of the pupil appears flat rather than round. You need to apply an initial 3000 nm of Z7 to get the astigmatism close.

02:12 - Image Analysis: 021313 -8272 0 0 0 0 0
02:19 - Image Analysis: 021921 -8457 0 0 0 0 0
Then John sent a further correction manually.
02:21 - Image Analysis: 022200 No GOOD pupils found... check input image!
Actuator still non-working. Images heart-shaped. Applied coma Z7 +1000 nm manually.
02:25 - Image Analysis: 022519 No GOOD pupils found... check input image! Try more coma (another Z7 +1000)
02:27 - Image Analysis: 022740 No GOOD pupils found... check input image! Go the other way Z7 -3000
02:29 - Image Analysis: 023011 Only Z4 correction but on the wrong pupil. Did not send. Applied Z7-500 Z4+1000
02:32 - Image Analysis: 023312 As above, got wrong pupils, Applied Z4 +2000 Z7 -500
02:35 - Image Analysis: 023626 -1424 2377 1392 591 2681 867 Sent
02:38 - Image Analysis: 023915 -688 -579 482 -369 -1229 -441

Standard Observations and Shutter Calibration

Seeing started as good as 0.9 to 1.2 arcsec, but then degraded to 2.7 arcsec.

02:42 - Preset to SA113 do the 8pos UsUB OB.
024604: seeing = 1.1". (SDT_Uspec)

Images 025458 and 025543 have either a jump, or horrible astigmatism. Mirror is probably having problems, probably unable to do photometry on these images. Do another IA after end of OB.

030850: seeing = 1.7 but due to the mirror funny behavior....
03:10 - Image Analysis: 031158 -1106 247 -92 -972 -873 40 Send
03:13 - Image Analysis: 031417 -88 -106 705 -76 -247 -50 Send

Enough, test the shutter.

03:16 - ShutterRamp 1 (from 0.1 to 3.0 s) (done @ airmass 1.18)
031744-033232 seeing around 1.3".

03:33 - ShutterRamp 2 (from 4.0 to 18.0 s)
033343-035349 pretty unstable seeing. Probably needs to be redone.

03:57 - Move to And XIV (V1) Image Analysis: 040036 926 335 -535 1053 -694 1000 Sent only Z4
Image Analysis: 040328 1802 218 -261 707 -317 1000 Z11 truncated. Sent only Z4 +900
Image Analysis: 040641 -726 347 140 1346 -552 1000 Z11 truncated, Sent only Z4 -726
Image Analysis: 041015 732 641 362 1199 299 1000

Seeing too bad for almost any program. If it stays like this, do Ryan's Asteroids but before take a standard field in all filters.

04:12 - Move to SA113, do the 4positions OBs (both Vgr and UsUB).
041540: seeing = 2.1"
041936: " = 1.9"

Seeing very variable (1.7-2.2) but overall horrible.

04:25 - Do the UsUB OB for SA113.
042635: seeing = 2.05"
043111: " = 2.0"

04:47 - Move to SA92 do the 4pos OBs First the UsUB
044804: seeing = 2.4"
045248: double stars: jump in the telescope? Mirror acting strange?

05:02 - SA92 4pos Vgr, and see what happens
050359: seeing = 1.8"
050757: " = 2.0"
051314: " = 2.7"

Issue 850 - Mirror Oscillations

No your eyes are not deceiving you....this is an October 2007 issue.

All evening we've been fighting the active optics and getting elongated images. We were trying to find telescope jumps, or wind or temperature gradients. Then we tried to blame actuator 140 for giving us excess astigmatism and funny shaped pupil boundaries.

At 22:00 MST we finally realized that the SX mirror was oscillating. It appears that the mirror is oscillating enough to elongate the image by several arcsec. The oscillation starts when the LBC asks PSF to move the mirror. (Sound familiar?) During oscillation, forces on the hardpoints were fluctuating up to 90 pounds+-. At 22:01, I was able to stop the oscillation by stopping and restarting the Z-servo loop for 2 sec.

During this period wind forces on the mirror were fluctuating below the 2# level on the hardpoints. Roughly 4 m/sec outside. And the telescope servo appeared stable. So it seems that the oscillation is caused by something in the cell and not by external excitation. Did we buy the 18 month warranty on those new seals?

Cell power supply voltages look normal at: 7.28, 15.37, -14.69, 19.69. Cell ambient temperature is 5 degC and slowly rising.

I was able to cure the problem for now by setting Outer Count to 10. Thanks for not removing that relic piece of software, Chris.

Probably we've wasted a total of 2 hours fighting the effects of this problem tonight and at the end of last night.

[Note added later: Retuned mirror cell Y servo loop to supress the problem the next day, but still haven't understood it.]

SDT Science Observations

05:14 - Move to Ryan Oct 15 deg epoch1
Doesn't make much sense to do IA, but let's try anyway.
Image Analysis: 051739 -185 -219 -740 2603 1240 1000 (not applied)

05:19 - Started Science OB [Oct 15 deg epoch 1]
052035: seeing = 3.0"
052558: " = 3.0"

05:30 - Go to [Oct 10 deg epoch 1] Image analysis: skipped
053315-053842

05:38 - Independent (by Ray Bertram with polaris monitor) measurement of seeing gives out 2.8".

05:45 - Move to [Oct 5 deg epoch 1] Image Analysis: skipped, seeing always around 3.0"
054621-055141

05:56 - Move to [Oct 0 deg epoch 1] Image analysis: skipped , seeing still around 2.7"
055751-060310

06:08 - Go to standard field SA92 , do the 4pos UsUB and Vgr OBs
061055-062134
062134: seeing ~ 2.0"

06:23 - SA92 Vgr
062339-063255
062856: seeing = 1.8"
063255: seeing = 1.7"

06:33 - Ryan's [Oct 15 deg epoch 2]
Image Analysis: 063732 -1492 -1317 -196 3022 1988 756 Sent
Image Analysis: 064058 -888 -45 -1310 -1442 -807 -413 Sent
Image Analysis: 064317 684 19 1811 -74 -393 -92 Sent
Image Analysis: 064505 -287 -263 403 -140 -173 -57 Sent

06:46 - Ryan [Oct 15 deg epoch 2]
064741: seeing = 1.75"
065315: " = 1.65"

06:58 - Ryan [Oct 10 deg epoch 2]
Image Analysis: 070005 246 589 -817 1284 716 932 Send
Image Analysis: 070209 -639 -310 -232 -632 -163 -363 Send

07:04 - Ryan [Oct 10 deg epoch 2]
070459: seeing = 1.9"
071026: " =

07:16 - Go to Ryan [Oct 5 deg epoch 2]
Image Analysis: 071824 343 237 -99 -283 -172 -322 Send

07:19 - [Oct 5 deg epoch 2]
072119: seeing = 1.5"
072643: seeing = 1.7"

07:31 - Go to Ryan [Oct deg 0 epoch 2]
Image analysis: 073300 -168 576 -139 -479 215 -93

07:34 - Sent OB [Oct 0 deg epoch 2]
073550: seeing = 1.7"
074108: " = 1.7"

07:48 - Go to Ryan [Oct deg 15 epoch 3]
Image analysis: 075055 -929 -927 68 221 130 -111 Send

07:53 - [Oct 15 deg epoch 3]
075330: seeing = 1.2"
075847: seeing =

08:04 - Go to Ryan [Oct 10 deg epoch 3]
Image analysis: 080706 932 339 -330 35 -388 -131

Go to target Oct 15 deg to redo the OB, too close to the zenith, images of 075847 show derotating problems typical of closeness to the zenith (85 degrees elevation).

08:09 - [Oct 15 deg epoch 3] (redo)
081214: seeing = 1.2"
081750: seeing = 1.3"

Pointing

Exposures taken near 85 degrees show evidence of rotation trailing. See 075847 crossing the meridian. But another exposure taken a few minutes later looks fine 081214 at 82 degrees. We turned on pointing logging near 08:05 UT so we can grab a pointing star near zenith.

082945 V5sec at EL=87.3 AZ=328 [1292,2858] Star needs to move toward upper left of the image, toward larger azimuth (according to WCS A) by 268 pixels. (But I'm not sure what that means relative to last night's measurements.)

SDT Science Observations Continue

08:30 - [Oct 10 deg epoch 3]
Image Analysis: 083449 623 151 106 -827 -505 13

08:38 - [Oct 10 deg epoch 3]
083918-084444
083918: seeing = 1.0"

08:50 - [Oct 5 deg epoch 3]
Image Analysis: 085231 177 502 -84 -95 -175 -149

08:54 - [Oct 5 deg epoch 3]
085456: seeing = 1.0"
090017: " = 1.0"

09:05 - [Oct 0 deg epoch 3]
Image Analysis: 090741 1018 -392 758 -6 30 256

09:08 - [Oct 0 deg epoch 3]
090956: seeing = 1.0"
091513: seeing = 0.9"

09:22 - Move to SXDS_1b (SDT_Uspec) airmass=1.30
Image Analysis: 092323 574 941 -1 -8 -113 74

09:24 - SXDS_1b
092758-094146 (OB aborted, seeing worse than specification)
092758: seeing = 1.4"
094146: seeing = 1.2"

09:35 - seeing rapidly degrading probably will abort the OB.

09:49 - aborted, move to AndXIV, V1
Image Analysis: 095516 -265 -331 -127 611 -200 0

09:57 - Launched AndXIV_V1
095758-100948
095758: seeing = 1.25
100948: " = 1.05

10:16 - go to AndXIV_B1
Image Analysis: 101715 -192 453 -69 326 20 115

10:19 - Launched AndXIV_B1 (airmass = 1.34)
102010-103154
102010: seeing = 1.1"
102600: " = 1.15"

10:38 - Go to AndXIV_V2 (airmass=1.43)
Launch directly w/out I.A., close enough to the previous one
103937-105129
103937: seeing = 1.1"
104537: " = 1.05"

10:57 - AndXIV_B2
Image Analysis: 105816 152 509 351 286 128 130

10:59 - Launched AndXIV_B2 (airmass = 1.54)
11:02 - Stopped OB because telescope lost tracking.....
11:01 - Jet crashes and reboots itself. Lost telescope tracking for 15 minutes.

11:10 - Problem fixed, IssueTrak filed, but AndXIV too low, moon rising, move to Stanek's M33.
Image Analysis: 111307 152 509 351 286 128 130

11:14 - Launched OB on M33
111628-

11:22 - Something went wrong with IA. Aborted OB, stars look horribly triangular.
Image Analysis: 112539 767 -41 -158 109 82 138

11:28 - Moon rising, high airmass, postpone Stanek's program and Move to Dwarf UMaII_1_V. No IA, the one just performed still valid
113235: seeing = 0.9"
113818:
114359:
114943: guiding stars showed some coma

0.9 arcsec images in V while observing UMaII at 1.4 airmasses!, but becoming comatic by the end of the 20 minute OB.

Standards

11:56 - go to SA98 4positions (UsUB)

12:15 - SA98 4positions (Vgr)

12:28 - go to SA95 4positions (Vgr)

12:38 - SA95 4positions (UsUB) aborted, too much light, go for morning flatfields

Morning Sky Flats

12:41 - SkyFlatTestG (exp.time = 1.0s) 124331: 7000 counts

12:44 - SkyFlatGrot90_5 (exp.time = 5.0s)
124519: 53932
124603: 63949
124648: 64870
lowered exp.time to 2.0s
124751: 39433
124837: 48073

12:49 - SkyFlatGrot270_5 (exp.time = 1.0s)
125028: 38399
125108: 45954
125148: 54647
125228: (saturated) 64839
125310: 64873

12:54 - SkyFlatTestU 125543: 15000

12:56 - SkyFlatU_5 (exp.time = 1.0s)
125649: 20201
125729: 23983
125816: 29054
125858: 34625
125938: 40557

13:00 - SKyFlatUrot90_5 (exp.time = 1.0s)
130101: 56285
130140: 64670 (saturated!)
130220: 64868 (saturated!)

Close Dome

13:03 - 25 biases

-- JohnHill - 08 Oct 2007
Topic revision: r7 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback